OpenStudy (anonymous):

What function best represents a sine function with an amplitude of 4, a period of π/6, and a midline at y = −2?

OpenStudy (crashonce):

ampltude means it will be like y=4sinx

OpenStudy (anonymous):

a. -2sin(x-π/6)+4 b. 4sin12x-2 c. 4sin(x-π/6)-2 d. -2sin12x+4

OpenStudy (anonymous):

Asin(Bx-C)+D where A is the amplitude, period is 2pi/B, D is the midline. so it would be 4sin(12x)-2 since the period is pi/6, pi/6 = 2pi/B, we can simplify this out to B=12. im 95% sure this is right. been a bit since ive done these.

OpenStudy (anonymous):

since there is no C (C is the phase shift), dont worry about that part.

OpenStudy (anonymous):

So it would definitely be b or c right? Because the amplitude is the first number?

OpenStudy (anonymous):

id say its definitely 4sin12x-2, but try to actually learn how it works first...

OpenStudy (anonymous):

Oh sorry I just saw that you put what you thought the answer was. Thank you so much!

OpenStudy (anonymous):

when solving for B - 2pi/B = pi/6 (the period), we can multiply B by both sides get Bpi/6=2pi, then we divide both sides by pi/6. we get B= 6x2pi/pi, the pi's cancel out and were left with B=12.

OpenStudy (anonymous):

when dealing with tan functions, the period is = pi/B rather then 2pi/B

OpenStudy (anonymous):

Yeah.. Thanks for helping me with this problem!